A standard gold bar stored at Fort Knox, Kentucky, is 7.00 inches long, 3.63 inches wide, and 1.75 inches tall. Gold has a density of 19,300 kg/m3. What is the mass of such a gold bar?
Answer #1
Concepts and reason

The concept required to solve this problem is density.

Initially, calculate the volume of the gold bar. Then, calculate the mass by using the equation of density.

Fundamentals

The volume of cuboid is,

V=lwtV = lwt

Here, ll is length, ww is width, and tt is height.

The density of a material is,

ρ=mV\rho = \frac{m}{V}

Here, mm is mass and VV is volume.

Substitute 7.00inches7.00{\rm{ inches}} for ll , 3.63inches3.63{\rm{ inches}} for ww , and 1.75inches1.75{\rm{ inches}} for tt in the equation V=lwtV = lwt .

V=(7.00inches)(3.63inches)(1.75inches)V = \left( {7.00{\rm{ inches}}} \right)\left( {3.63{\rm{ inches}}} \right)\left( {1.75{\rm{ inches}}} \right)

Convert inches to meter by multiplying with (1m39.37inches)\left( {\frac{{1{\rm{ m}}}}{{39.37{\rm{ inches}}}}} \right) .

V=((7.00inches(1m39.37inches))(3.63inches(1m39.37inches))(1.75inches(1m39.37inches)))=7.29×10−4m3\begin{array}{c}\\V = \left( \begin{array}{l}\\\left( {7.00{\rm{ inches}}\left( {\frac{{1{\rm{ m}}}}{{39.37{\rm{ inches}}}}} \right)} \right)\left( {3.63{\rm{ inches}}\left( {\frac{{1{\rm{ m}}}}{{39.37{\rm{ inches}}}}} \right)} \right)\\\\\left( {1.75{\rm{ inches}}\left( {\frac{{1{\rm{ m}}}}{{39.37{\rm{ inches}}}}} \right)} \right)\\\end{array} \right)\\\\ = 7.29 \times {10^{ - 4}}{\rm{ }}{{\rm{m}}^3}\\\end{array}

Rearrange the density equation to solve for mass.

ρ=mVm=ρV\begin{array}{c}\\\rho = \frac{m}{V}\\\\m = \rho V\\\end{array}

Substitute 19300kg/m319300{\rm{ kg/}}{{\rm{m}}^3} for ρ\rho and 7.29×10−4m37.29 \times {10^{ - 4}}{\rm{ }}{{\rm{m}}^3} for VV in the equation m=ρVm = \rho V .

m=(19300kg/m3)(7.29×10−4m3)=14.07kg\begin{array}{c}\\m = \left( {19300{\rm{ kg/}}{{\rm{m}}^3}} \right)\left( {7.29 \times {{10}^{ - 4}}{\rm{ }}{{\rm{m}}^3}} \right)\\\\ = 14.07{\rm{ kg}}\\\end{array}

Ans:

The mass of the gold bar is 14.07kg14.07{\rm{ kg}} .